OpenStudy (kendricklamar2014):

Two similar cylindrical vases have diameters of 6 in. and 8 in. The smaller vase has a surface area of 1,960 in^2 and a volume of 339 in^3. Find the surface area and the volume of the larger vase.

OpenStudy (kendricklamar2014):

@aaronq, @AccessDenied, @HelpBlahBlahBlah, @hhelpplzzzz, @mathmale, @thomaster can you help me

OpenStudy (mathmale):

I think it best to assume that the vases have bottoms (which are circular), and that the two vases have the same heights but different diameters. What are those diameters? What are the corresponding radii? What is the formula for the circumference of a circle? Once you have this info, we can move forward to tackle the rest of this problem.

OpenStudy (kendricklamar2014):

How do you find the diameters?

OpenStudy (anonymous):

As they are similar ,means they have same height. for smaller one,r=3 volume=339 vol=pi*r^2*h; r is radius, h is height or, pi*(3^2)*h=339 h=12 for larger one,r=4 vol=pi*(4^2)*12 vol=602.88 in.3 Surface area=2*pi*r*h+2*pi*r^2 =2*pi*(4*12+4^2) =401.92 in.2

OpenStudy (kendricklamar2014):

OK, thank you :)